### v1

# FaceShield[Uncaptioned image]: Explainable Face Anti-Spoofing with Multimodal Large Language Models

FaceShield[Uncaptioned image]: Explainable Face Anti-Spoofing with Multimodal Large Language Models
Hongyang Wang1, Yichen Shi2, Zhuofu Tao3, Yuhao Gao1, Liepiao Zhang4, Xun Lin5
Jun Feng1, Xiaochen Yuan6, Zitong Yu5🖂, Xiaochun Cao7
1Shijiazhuang Tiedao University   2Shanghai Jiao Tong University  3UCLA  4GRGBanking
5Great Bay University   6Macao Polytechnic University   7Shenzhen Campus of Sun Yat-sen University
Equal contribution      🖂 Corresponding author
Figure 1: FaceShield Multi-task Response Demonstration. This figure illustrates the model’s performance across four tasks: coarse-grained classification, fine-grained classification, reasoning, and localization. Coarse-grained classification distinguishes real from spoofed faces, while fine-grained classification identifies specific attack types, such as print attacks. The reasoning task provides explanations for spoofing by analyzing features like lighting and symmetry. Localization precisely detects spoofed regions. This illustration highlights FaceShield’s capability to handle a range of complex questions with accuracy.
Face anti-spoofing (FAS) is crucial for protecting facial recognition systems from presentation attacks. Previous methods approached this task as a classification problem, lacking interpretability and reasoning behind the predicted results. Recently, multimodal large language models (MLLMs) have shown strong capabilities in perception, reasoning, and decision-making in visual tasks. However, there is currently no universal and comprehensive MLLM and dataset specifically designed for FAS task. To address this gap, we propose FaceShield, a MLLM for FAS, along with the corresponding pre-training and supervised fine-tuning (SFT) datasets, FaceShield-pre10K and FaceShield-sft45K. FaceShield is capable of determining the authenticity of faces, identifying types of spoofing attacks, providing reasoning for its judgments, and detecting attack areas. Specifically, we employ spoof-aware vision perception (SAVP) that incorporates both the original image and auxiliary information based on prior knowledge. We then use an prompt-guided vision token masking (PVTM) strategy to random mask vision tokens, thereby improving the model’s generalization ability. We conducted extensive experiments on three benchmark datasets, demonstrating that FaceShield significantly outperforms previous deep learning models and general MLLMs on four FAS tasks, i.e., coarse-grained classification, fine-grained classification, reasoning, and attack localization. Our instruction datasets, protocols, and codes will be released soon.
1 Introduction
Face anti-spoofing (FAS) is essential in facial recognition systems, ensuring that presentation attacks (PAs), such as print, replay, and 3D wearable masks, are effectively prevented. It has attracted considerable interest in industry and academia in the past decade.
Existing deep learning FAS models can be categorized into two types: vision-based methods and vision-language-based methods. As shown in Fig. 2(a), vision-based methods rely solely on image data (e.g., RGB, Depth, Infrared(IR)) and binary labels to train CNNs [66, 67] or ViTs [15, 70] for FAS. While they can achieve satisfactory results against known attack types and environments, these methods are prone to overfitting on spurious correlations and lack strong extrapolation capabilities. As illustrated in Fig. 2(b), Vision-language-based methods do not use binary labels but instead train CLIPs with image-text pairs [55, 36, 43]. The text labels in these methods provide more domain-agnostic information, enhancing models’ generalization capability. Although these FAS models demonstrate some recognition capabilities, they still face challenges such as limited generalization ability, poor interpretability, and a lack of capability for fine-grained localization of attack regions.
Recently, MLLMs have shown remarkable capabilities across various visual tasks, such as remote sensing [76, 29, 44], medical imaging [32, 57], and deepfake detection  [64, 25]. By leveraging the general capabilities of language foundation models alongside the visual information extracted by vision towers, these specialized MLLMs integrate perception, reasoning, and decision-making within a single model. Regarding the FAS task, SHIELD [53] conducted extensive evaluations on existing general-purpose MLLMs, revealing that their performance on FAS tasks still has room for improvement. [75] introduced a model capable of performing classification and description attack type. However, the model is limited in its ability to handle more nuanced tasks, such as identifying specific attack types, reasoning, and localizing spoofed areas. These limitations underscore the broader challenges in training FAS MLLMs, including: (1) a lack of pretraining and SFT datasets specific to FAS tasks, (2) the need to extend traditional FAS tasks to fully exploit MLLM capabilities, and (3) the difficulty for general-purpose vision towers to capture the subtle distinctions between real faces and PAs, unlike with natural images.
Motivated by the above discussion, in this paper, we expand the traditional FAS task to include four sub-tasks (see Fig. 1 for examples): coarse-grained classification, fine-grained classification, reasoning, and attack localization. We then introduce FaceShield, an MLLM specifically designed for these tasks. As can be seen from Fig. 2(c), we propose a pretraining and SFT dataset generation pipeline. This pipeline constructs two multimodal FAS instruction datasets containing 50k dialogues for FaceShield training. To the best of our knowledge, FaceShield is the first FAS MLLM, equipped with multiple detection capabilities. Additionally, FaceShield-pre10K and FaceShield-sft45K are the first high-quality datasets that can be used to train a FAS-specific MLLM. Our main contributions include:
  • We develop a novel data generation pipeline that utilizes a MLLM and predefined prompts, and construct two multimodal FAS instruction datasets (i.e., FaceShield-pre10K and FaceShield-sft45K) with 12 attack types. To our best knowledge, these are the first multitask instruction datasets for the FAS community.
  • We propose FaceShield, the first multitask MLLM for FAS that is capable of coarse-grained classification, fine-grained classification, reasoning, and attack localization. FaceShield  utilizes the Spoof-Aware Vision Perception (SAVP) and Prompt-guided Visual Token Masking (PVTM) strategies to enhance the discrimination of confusing attack areas.
  • Extensive experiments demonstrate that FaceShield significantly outperforms previous specialized FAS models and general MLLMs across multiple datasets in various FAS evaluation tasks.
Figure 2: Pipelines of different FAS methods (a) traditional deep learning models, (b) multimodal models, and (c) MLLM
2 Related Work
Figure 3: Construction pipeline of our proposed instruction datasets (i.e., FaceShield-pre10K and FaceShield-sft45K).The initial datasets (WMCA, PADISI, SiW-Mv2) are combined to form a uni-class dataset covering 12 spoofing types, with selected images annotated for visual grounding. Using MLLM with structured prompts, we generate two datasets: a pretraining dataset and an SFT dataset divided into four tasks (coarse-grained classification, fine-grained classification, reasoning, and localization). The pretraining data is filtered by CLIP for similarity, producing the FaceShield-pre10k dataset. SFT data undergoes multi-level filtering (LLM-based, keywords, and human reviews), followed by augmentation, resulting in the FaceShield-sft45k dataset.
Face Anti-Spoofing. Previous works usually employed convolutional neural networks (CNNs) [66, 67] or vision transformers (ViTs) [15, 70, 10] as backbone networks, integrating auxiliary priors for live/spoof classification (e.g., reflection light [6], depth reconstruction [6, 41, 3], rPPG signals [33, 65, 41, 73], and gradient information [68, 8]) to detect PAs. While these methods perform well in source domains, their performance degrades significantly in unseen domains due to domain shifts and unseen attacks [71]. Recently, many domain-generalized FAS methods [71] have been proposed to address these cross-domain challenges. Adversarial training [27, 26] and feature disentanglement techniques [42, 28, 17, 62] have been explored to decouple image features into domain-invariant (content) and domain-specific (style) features, preventing the model from learning spurious correlations between domain-specific features and labels. Some approaches employ meta-learning [48, 52, 7], data augmentation [9], or data synthesis methods [14, 39] to simulate domain shifts, enabling the model to learn a broader data distribution during training. Others perform domain [26, 58, 21, 60, 78] or gradient [30] alignment to constrain the relative relationships between sample embeddings across different source domains, aiming to learn more generalized decision boundaries. More recently, some methods have adopted vision-language models (VLMs), particularly CLIP [49], to detect PAs using text guidance. By leveraging the reasoning capabilities acquired during large-scale pretraining, VLM-based approaches [56, 36, 61, 37, 40] have shown to improve the generalization of FAS. Unlike traditional binary live/spoof classification, VLM-based methods classify images by computing similarities between visual features and textual descriptions of live and spoof faces, leveraging more causal information [36].
Existing FAS methods lack interpretable reasoning capabilities, e.g., language-based explanations, to distinguish between live and spoof faces. A concurrent work [75] collect 12 datasets and described images based on attack types and keywords. Using instruction tuning, they develop a MLLM capable of classifying and providing detailed descriptions of attacks. Meanwhile, these approaches fail to integrate classification, localization, and reasoning tasks simultaneously, limiting their adaptability and reliability in real-world applications, where versatile understanding is critical for reliable decision-making. Additionally, existing FAS datasets [61, 37] lack annotations for reasoning processes, further hindering the development of reasoning-capable FAS methods.
Figure 4: Comparison of different model architectures. (a) Single modal baseline model. (b) Proposed model without Prompt-Guided Vision Token Masking (PVTM). (c) Proposed model with PVTM. The input image is first encoded by the vision encoder, and PVTM partitions the encoded features into RGB, SAV, and Masked tokens. These visual tokens are passed to the projector, while the prompt is tokenized separately, producing text tokens. The combined visual and text tokens are then fed into the Large Language Model (LLM) with LoRA, producing the final output.
Multimodal Large Language Model. MLLMs, such as GPT-4V [2], LLaVA [38], Bunny [20] have demonstrated successful performances in general domains [13, 31, 24], attracting considerable attention from researchers. Researchers have developed task-specific MLLMs in various fields under vision tasks, including remote sensing [76, 29, 44], medical [32, 57], aesthetic evaluation [23], deepfake detection [64, 25] and Visual Grounding [4, 63, 46]. Unlike traditional models, MLLMs can handle multiple tasks within a single model and exhibit strong performance in unknown domains. FakeShield [64] proposes an explainable multi-task model for image forgery detection and localization, capable of making judgments based on tampering clues at both the image and pixel levels. The language foundation model contains scarce knowledge of face attacks, resulting in poor performance of existing MLLMs on FAS tasks. Our constructed datasets can introduce abundant FAS-related knowledge to MLLM, which through SAVP further enhances its ability to distinguish between real faces and PAs, and PVTM enhances the generalizability of the extracted features.
3 FaceShield-pre10K and FaceShield-sft45K
3.1 Dataset Collection
Fig. 3(a) shows the collection and localization coordinate annotation process of existing FAS datasets. Based on class types of WMCA (W) [16], PADSIS (P) [51] and SiW-Mv2 (S) [18] datasets, we organize all annotation categories into 12 distinct types: Bonafide, Fakehead, Print, Glasses, Replay, Paper mask, Flexible mask, Rigid mask, Partial Eye, Partial Mouth, Makeup, and Tattoo. We have also re-annotated the original dataset images at both the image and region levels. Finally, we obtained 12,091 images with class labels and 3,139 images with bounding box labels. More details and examples can be found in our Appendix.
3.2 Instruction Construction
After collecting images, class labels, and attack regions, as shown in Fig. 3(b), we use Bunny-Llama-3-8B-V [20] to assist in generating the pretrain dataset and the SFT dataset. Specifically, we first use a system prompt as a preamble to enhance the performance of responses from the MLLM assistant. Class labels are embedded in the prompt, providing descriptions of some image content. Subsequently, given the task type, the MLLM assistant generates multiple question-and-answer pairs for the task, following the format provided in the few-shot examples.
For the construction of the pretraining dataset FaceShield-pre10K, we only use the MLLM assistant to generate descriptions of the images, without any other information. The image and its corresponding generated description are filtered using a CLIP [50] score, removing image-text pairs with a similarity of less than 15%.
For the task-specific instruction dataset FaceShield-sft45K, after generating question-answer (QA) pairs using MLLM, manual filtering, and specific keyword filtering are applied to remove incorrect QA pairs. We enhance the diversity of the content in questions and answers using Llama3 [12] on the golden seed dataset obtained after filtering, to improve the dialogue capabilities of the model trained on this dataset. The specific content and responses for each task are as follows:
  • Coarse-grained classification: MLLM directly responds with labels for face images, identifying them as either real faces or PAs.
  • Fine-grained classification: The responses from the MLLM include not only real or PAs but also judgments on the specific categories of the PAs.
  • Reasoning: MLLM provides descriptions and reasoning from multiple perspectives on the image and delivers the final judgment.
  • Attack localization: MLLM first evaluates the faces in the image, if PA is detected, it provides the specific coordinates of the attack area.
4 FaceShield
Our goal is to train a FAS task-specific MLLM with two main objectives: 1) Enhance the visual encoder’s ability to extract features from real faces and presentation attacks, and 2) Utilize the extensive knowledge stored in the LLM to improve the model’s generalization capabilities when facing unknown domains. A naive training approach is shown in Fig.  4(a), where direct pre-training and SFT are conducted using RGB images and constructed QA data. However, the high similarity between real faces and PAs in RGB appearance poses significant challenges to this method. As shown in Fig. 4(b), Spoof-Aware Vision Perception (SAVP) combines images preprocessed based on prior knowledge, by extracting predefined local descriptor operators [72], with the original RGB image. Our complete model framework is shown in Fig. 4(c), RGB images and the extracted local descriptor images are fed into the vision encoder to extract vision token $V_{RGB}$ and $V_{SAV}$, respectively. These features are then processed through the Prompt-Guided Vision Token Masking (PVTM) module, which extracts highly generalizable vision tokens. These tokens are sent to a projector to align with text prompt token $P$ to produce $V_{align}$, which is then fed into the language foundation model for inference result $\mathcal{Y}$ as follows:
$V_{align}=\text{Projection}(V_{RGB},V_{SAV})$ (1)
$\mathcal{Y}=\text{MLLM}(V_{align},P)$ (2)
4.1 Spoof-Aware Vision Perception
As shown in Fig. 4(a), bonafide faces and PAs lack distinct discriminative features in RGB-based appearance space, whereas local descriptors [66, 72] extracted through image preprocessing can enhance their subtle live/spoof clues. As shown in Fig. 4(b), we extract features from the original images using Local Binary Pattern (LBP) [47], Gray, and Histogram of Oriented Gradients (HOG) [11], and concatenate them. LBP and Gray-specific computations are as follows:
$LBP$ $=\sum_{i=0}^{P-1}s(g_{i}-g_{c})\cdot 2^{i},$ (3)
$s(x)$ =$=\begin{cases}1&\text{if }x\geq 0,\\ 0&\text{otherwise},\end{cases}$
where $g_{c}$ is the pixel value of the central pixel in the considered neighborhood, and $g_{i}$ represents the pixel values of the $P$ surrounding pixels.
$\text{Gray}(I)=0.299\cdot R+0.587\cdot G+0.114\cdot B,$ (4)
where $R$, $G$, and $B$ are the red, green, and blue intensity values of the pixel, respectively.
The HOG calculates the gradient magnitude and direction at each pixel using edge detection operators and then divides the image into small, overlapping cells. Within each cell, gradients are binned according to their direction into histograms. Histograms from the cells within each block are concatenated and normalized based on the block’s overall gradient energy. The final HOG descriptor is formed by the vector of these normalized histograms from all blocks.
We perform the above three steps of feature extraction on the original image, then concatenate these as three channels to form a complete image. This composite image is then fed into the vision encoder to extract features as complementary information. The spoof-aware vision token $V_{SAV}$ and final vision input $V$ for FaceShield as follows:
$V_{SAV}=\text{Encoder}(\text{Concat}[LBP,Gray,HOG])$ (5)
$V=\text{Concat}[V_{RGB},V_{SAV}])$ (6)
Figure 5: Prompt-guided Vision Token Masking Module (PVTM)
4.2 Prompt-Guided Vision Token Masking
To further enhance the alignment between visual features and text prompts, and alleviate overfitting on spurious correlations, we leverage text prompts to guide token selection after the visual encoder. As shown in Fig. 5, the text tokens extracted from the prompt are pooled [54], and then their similarity with all visual tokens is calculated. We assume that visual tokens with higher similarity are more relevant to subsequent tasks and even for the same image, the important visual tokens may not be consistent across different tasks. The calculation of similarity between visual tokens $V_{i}$ and text prompt tokens $P$ is as follows:
$Sim(V_{i},P)=\frac{V_{i}\cdot P}{\|V_{i}\|\|P\|}$ (7)
Subsequently, we apply a softmax function to the similarities between all $V_{i}$ and $P$, using the resulting values $S^{i}_{rank}$ as an importance metric for each visual token. We then rank all tokens based on this importance calculated as follows:
$S^{i}_{rank}(V_{i},P)=\frac{e^{S(V_{i},P)}}{\sum_{j}e^{S(V_{j},P)}}$ (8)
Afterward, we retain the top $k\%$ of vision tokens in importance. The remaining tokens are then randomly masked with a probability of $p\%$, reducing the influence of less important tokens while keeping acceptable information loss for the final decision-making process.
4.3 Training Details
We employ a two-stage training strategy. The first stage is the pre-training stage, where we perform continual pretrain of existing weights. In this stage, we align the visual embeddings from a pre-trained vision encoder with the text embeddings using FaceShield-pre10K. The second stage is the supervised fine-tuning (SFT) stage, where we apply visual instruction tuning to fully leverage the MLLM’s capabilities across various multimodal tasks within our specific domain using FaceShield-sft45K. For LLM tuning, we incorporate LoRA [22] and set the parameters with a hidden dimension of 128 and a scaling factor of 256 in every transformer block. In both stages, we employ cross-entropy loss for next-token prediction. During the pre-training stage, only the projector after the vision encoder and PVTM are optimized for one epoch. In the SFT stage, we fine-tune the LoRA layers in LLM as well as the projector.
Table 1: Intra-dataset results on coarse-grained classification.
Method ACC(%) $\uparrow$ HTER(%) $\downarrow$
Traditional
ResNet [19] 97.55 2.32
PatchNet [59] 98.22 1.78
CoOp [77] 98.73 1.27
MLLM
LLaVA [38] 65.54 27.76
Qwen-VL [4] 51.94 38.70
Minigpt4 [79] 26.86 65.50
Bunny [20] 81.20 17.87
Bunny (fine-tuned) [20] 98.23 1.52
FaceShield (Ours) 99.41 ± 0.06 0.53 ± 0.06
5 Experiments and Results
5.1 Protocols and Evaluation Metrics
We select 10% of each source dataset in FaceShield-sft45K to form three test subsets (W, S, P). For the coarse-grained classification and fine-grained classificationtask, we conduct both intra-dataset and cross-dataset testing. For the reasoning, and attack localization tasks, we employ intra-dataset evaluation. The intra-dataset test uses all data from the source datasets for training and evaluation on the combined test sets (W&S&P). For cross-dataset testing, we select two datasets for pre-training and SFT, and then test using the test set from the remaining dataset. For example, we might train on W and S, and test using the test set from P. For the coarse-grained classification, fine-grained classification, and reasoning tasks, we use Half Total Error Rate (HTER) [69] or accuracy (ACC) as evaluation metrics. Additionally, we evaluate the reasoning process for the reasoning task using BLEU [45], ROUGE-L [34], and METEOR [5], and evaluate the attack localizationtask using AP@40 and AP@50.
5.2 Implementation Details
We use Siglip [74] as the visual encoder and Phi-3 [1] as the language foundation model. PVTM retains the top 10% of the most important tokens and randomly masks 5% of the tokens in the remaining 90%. Adam optimizer is used in the pretrain stage with a learning rate of $5\times 10^{-4}$. As for SFT stage, we decrease the learning rate to $2\times 10^{-4}$. All experiments are conducted on a single NVIDIA A100 GPU. Each experiment is repeated 10 times on the model, and the final results are reported as the mean ± standard deviation.
Table 2: Cross-dataset results on coarse-grained classification. $W$, $S$, and $P$ denote WMCA, SiW-Mv2, and PADISI, respectively.
Method ACC(%) $\uparrow$ HTER(%) $\downarrow$
W & S P
ResNet [19] 46.12 50.00
PatchNet [59] 77.18 22.87
IADG [78] 72.96 27.01
FaceShield (Ours) 93.17 ± 0.22 6.37 ± 0.21
W & P S
ResNet [19] 53.36 49.16
PatchNet [59] 56.16 45.37
IADG [78] 57.20 42.81
FaceShield (Ours) 89.93 ± 0.15 10.3 ± 0.14
S & P W
ResNet [19] 74.01 29.75
PatchNet [59] 78.15 41.50
IADG [78] 78.55 26.27
FaceShield (Ours) 92.56 ± 0.08 5.71 ± 0.08
Table 3: Results of fine-grained classification task.
Method ACC(%) $\uparrow$
LLaVA [38] 16.39
Qwen-VL [4] 16.55
Minigpt4 [79] 19.51
Bunny [20] 27.03
Bunny (Fine-tuned) [20] 94.43
FaceShield (Ours) 95.81 ± 0.11
Table 4: Results of the reasoning task with metrics BLEU, ROUGE-L, METEOR, ACC, and HTER.
Method BLEU-1 (%) $\uparrow$ BLEU-2 (%) $\uparrow$ BLEU-3 (%) $\uparrow$ BLEU-4 (%) $\uparrow$ ROUGE-L (%) $\uparrow$ METEOR (%) $\uparrow$ ACC (%) $\uparrow$ HTER (%) $\downarrow$
LLaVA [38] 45.05 31.75 23.42 17.80 30.51 25.52 37.84 50.11
Minigpt4 [79] 17.85 7.85 3.53 1.94 27.54 21.88 33.86 50.00
Qwen-VL [4] 20.92 14.53 11.01 8.77 21.45 12.49 47.64 41.49
Bunny [20] 33.64 27.12 22.65 19.33 36.74 19.12 50.68 39.73
Bunny(fine-tuned) [20] 89.57 86.96 84.91 81.29 80.15 51.64 98.56 1.16
FaceShield (Ours) 90.89 ± 0.14 88.02 ± 0.15 85.75 ± 0.17 83.98 ± 0.19 82.98 ± 0.20 53.10 ± 0.16 99.29 ± 0.04 0.57 ± 0.04
5.3 Comparison with Existing Methods
5.3.1 Coarse-Grained Classification Task
For the coarse-grained classification task, we compare FaceShield with state-of-the-art FAS methods  [19, 59, 77, 78] and open-source MLLM models [38, 4, 79, 20].
Intra-dataset Testing. Table 1 demonstrates that Our FaceShield significantly outperforms three representative traditional FAS methods [19, 59, 77] in within-domain experiments. Moreover, our performance greatly exceeds the zero-shot capabilities of general MLLM. We also fine-tune the open-source MLLM (i.e., Bunny [20]), selecting RGB images and language data from the dataset to conduct experiments on Bunny. We find that FaceShield also surpasses the well-tuned MLLM (Bunny) with 1% HTER decrease.
Cross-dataset Testing. Table 2 shows the performance of FaceShield in cross-domain scenarios, where we trained on two out of three selected datasets and tested on one. FaceShield demonstrates performance far exceeding traditional FAS models in cross-domain scenarios. Under the S&PW protocol, it achieves the HTER of 5.72%, showcasing FaceShield’s strong generalization capabilities compared to traditional methods.
5.3.2 Fine-Grained Classification Task
Table 3 shows the results under the fine-grained classification task. For open-source MLLMs [38, 4, 79, 20], we incorporated 12 types of attacks into the prompt, allowing it to respond with the correct type. For the fine-tuned MLLM and our FaceShield, we selected keywords from the responses for evaluation. It is evident that supervised fine-tuning can significantly improve the model’s performance, with FaceShieldachieving the best results. Cross-dataset results can be found in the Appendix due to space constraints.
5.3.3 Reasoning Task
We also explore the models’ reasoning capacity and Table 4 displays the results for the reasoning task. General MLLMs perform poorly in both classification and reasoning. FaceShield significantly outperforms general MLLMs in both the reasoning process and judgment, and also exceeds open-source MLLMs (e.g., Bunny) fine-tuned on our dataset. FaceShield not only provides accurate results but also delivers detailed and correct reasoning, effectively enhancing the explainability of FAS methods.
5.3.4 Attack Localization Task
It is vital to locate the spoof regions for explainable FAS, and Table 5 presents the results for the attack localization task. Due to the scarcity of attack localization annotations in general pre-trained datasets, general MLLMs perform poorly on this task. In contrast, FaceShield shows excellent results, achieving over 95% for both AP@40 and AP@50. It accurately locates attack areas, providing new insights into attack region detection for FAS tasks.
Table 5: Results of the attack localization task with metrics AP@40 and AP@50.
Method AP@40 (%) $\uparrow$ AP@50 (%) $\uparrow$
Qwen-VL [4] 2.07 1.49
Lenna [63] 37.77 35.41
Sphinx [35] 47.86 46.30
Bunny [20] 73.50 71.65
Bunny(fine-tuned) [20] 92.30 89.71
FaceShield (Ours) 97.78 ± 0.21 95.60 ± 0.19
Table 6: Ablation results across different tasks.
SAVP PVTM Coarse-grained classification Fine-grained classification Reasoning Attack localization
ACC(%) $\uparrow$ HTER(%) $\downarrow$ ACC(%) $\uparrow$ ACC(%) $\uparrow$ HTER(%) $\downarrow$ AP@40(%) $\uparrow$ AP@50(%) $\uparrow$
$\times$ $\times$ 98.23 1.52 94.43 98.56 1.16 92.30 89.71
$\checkmark$ $\times$ 98.73 1.06 94.59 99.41 0.48 97.09 95.21
$\checkmark$ $\checkmark$ $\mathbf{99.41\pm 0.06}$ $\mathbf{0.53\pm 0.06}$ $\mathbf{95.81\pm 0.11}$ $99.29\pm 0.04$ $0.57\pm 0.04$ $\mathbf{97.78\pm 0.21}$ $\mathbf{95.6\pm 0.19}$
Figure 6: Comparison of performance after fine-tuning using our proposed dataset on LLaVA [38] and Bunny [20] models.
5.4 Ablation Study
Effectiveness of the proposed datasets. We conduct pre-training and SFT with our proposed FaceShield-pre10K and FaceShield-sft45K on LLaVA [38] and Bunny [20] to evaluate the efficacy and generalization of our instruction tuning datasets. The results in Fig. 6 demonstrate that fine-tuning on our dataset significantly enhances performance across three tasks for both LLaVA and Bunny. This demonstrates that our dataset can effectively enrich general MLLMs with substantial FAS-related knowledge and boost their capabilities in FAS tasks. It also validates the efficacy of our advanced dataset construction pipeline.
Figure 7: Ablation of visual token masking ratio $p$ in PVTM on (a) three tasks (i.e., coarse- & fine-grained classification, and reasoning), and (b) attack localization task.
Effectiveness of SAVP. Results from the first two rows in Table 6 show that leveraging local descriptors as complementary visual inputs significantly improves performance across four tasks, particularly in the attack localization task, where AP@40 and AP@50 increased by 5.6% and 5.94%, respectively. It indicates that prior knowledge-based auxiliary information significantly enhances the model’s ability to distinguish easily confusable facial images. The local live/spoof details within the auxiliary data proves especially valuable for fine-grained attack region detection tasks.
Figure 8: Ablation of visual token retain ratio $k$ in PVTM on (a) three tasks (i.e., coarse- & fine-grained classification, and reasoning), and (b) attack localization task.
Effectiveness of PVTM. It can be seen from the last two rows of Table 6 that the proposed PVTM provides reasonable improvements for FaceShield across three (coarse- and fine-grained classification, and attack localization) tasks. It indicates that masking less important tokens helps prevent the model from task-unrelated noises and spurious correlations. However, PVTM leads to a slight performance decrease on the reasoning task. This might be because masking partial visual tokens may compromise overall image perception and lose information for reasoning.
We further conduct studies of PVTM with different visual token masking ratios $p$, as shown in Fig. 7(a) and (b). Through our experiments for three (coarse- and fine-grained classification, and attack localization) tasks, we find that masking 5% of the tokens strikes an optimal balance between reducing spurious correlations and preserving essential information. However, we cannot find any improvement when masking more tokens due to severe information loss. Besides, we also investigate varying proportions of visual token retain ratio $k$, as illustrated in Fig. 8(a) and (b). We preserve 0%, 10%, 20%, and 30% of the tokens, respectively, and then randomly mask $p$=5the remaining tokens. The results show that retaining $k$=10visual tokens with strong importance achieves optimal performance. However, the more tokens are preserved, the poorer the model performs, suggesting that as the importance of tokens decreases, the likelihood of spurious correlations increases.
Figure 9: Importance visualization of (a) RGB vision tokens for classification tasks and (b) SAV tokens for attack localization task.
6 Visualization and Analysis
Here, we visualize the importance of RGB and SAV tokens. Fig. 9(a) displays the significance of visual tokens in relation to the prompt text for images from three different categories. For the coarse- and fine-grained classification tasks, it is evident that the tokens sensitive to prompt are mainly clustered in the facial regions. The features of these areas play a decisive role in determining the final outcome. Fig. 9(b) demonstrates how SAV tokens assist in the attack localization task. For this task, RGB tokens sometimes fail to provide accurate localization information. As shown in the left of Fig. 9(b), for the eye mask region, the importance visualization of RGB tokens is quite dispersed. However, in the SAV token visualization, the significance of the eye attack region is clearly highlighted. The right part of Fig. 9(b) shows how RGB and SAV tokens together provide supportive information for the attack localization task, including the eye mask attack area and the holding hands.
7 Conclusion
In this paper, we expand the FAS task into four sub-tasks: coarse-grained classification, fine-grained classification, reasoning, and attack localization, and propose the FaceShield, a specialized MLLM tailored for these FAS tasks. Considering the specific training data requirements of MLLM, we establish a pipeline for constructing datasets tailored for FAS task pre-training and supervised fine-tuning, resulting in the creation of the FaceShield-pre10K and FaceShield-sft45K datasets. This paper represents a preliminary exploration of MLLM for FAS task, and future works will focus on incorporating multiple visual modalities and refining the granularity of attack region localization.